OpenStudy (anonymous):

2t+8-t=-3 Please solve the equation!!

OpenStudy (anonymous):

2t+8-t=-3 2t-t=-3-8 t=-11 that's the value of "t". Check: 2(-11)+8-(-11)=-3 -22+8+11=-3 -14+11=-3 -3=-3 Correct. Then "-3" is your answer. :)